OpenStudy (anonymous):

A ball is thrown upward with a speed of 40 feet per second from the edge of a cliff 500 feet above the ground. What is the speed of the ball when it hits the ground? Use acceleration due to gravity as –32 feet per second squared and approximate your answer to 3 decimal places.

OpenStudy (anonymous):

@aaronq @dan815 @freckles @IrishBoy123 @Preetha @zepdrix @TheSmartOne

OpenStudy (irishboy123):

@familyguymath the classic 3 equations that are worth remembering - always- are \[v^2 = u^2 + 2 a x\] \[x = ut + \frac{1}{2}a t^2\] \[v = u + at\] listed in reverse [of normal] order. the first is oft-forgotten yet so powerful. but it is a most direct consequence of basic energy conservation: \[mgx = \frac{1}{2} mv^2 \]!! ciao

OpenStudy (anonymous):

@IrishBoy123 this doesnt help me w/ any of the info i was given

OpenStudy (irishboy123):

@familyguymath it does. just look at the first formula or consider the energy equation. either way, you have all the information you need.

OpenStudy (anonymous):

@IrishBoy123